# Mantis: A Simulation-Grounded Foundation Model for Disease Forecasting

### v3

@@ Discussion
Mantis demonstrates that accurate forecasting across diverse diseases can be achieved through training on simulated outbreaks, without access to real-world data. By learning from mechanistic simulations, Mantis delivers competitive performance across real-world forecasting tasks spanning multiple transmission modes, outbreak phases, and surveillance systems. These capabilities position Mantis as a practical tool for public health use, especially in low-resource or emerging outbreak settings where data may be limited or delayed. Because Mantis can generate forecasts with only a short initial observation window, it enables early situational awareness when conventional models require longer histories to achieve comparable accuracy. Mantis’s advantage in data-scarce settings stems from its ability to avoid overfitting to sparse initial data. Rather than extrapolating confidently from a handful of observations, Mantis produces appropriately hedged forecasts that reflect the true uncertainty across potential outbreak trajectories, enabling more reliable probabilistic planning even when historical context is minimal.
@@
The implications extend beyond public health. Simulation-supervised learning allows mechanistic models to embed scientific knowledge into neural networks, enabling them to generalize beyond observed data [8, 9]. This paradigm offers a potential solution to a fundamental challenge in epidemiological modeling: real-world disease data are limited by privacy concerns, reporting inconsistencies, and the infrequency of major outbreaks. Simulations, by contrast, can generate unlimited training data spanning diverse scenarios. Whether simulation-based training ultimately proves superior to training on massive real-world datasets—if such data could be assembled—remains an open question. However, given current data availability constraints, the simulation-based approach offers a practical path forward for building general-purpose forecasting systems.
@@
Mantis represents an initial attempt at a foundation model for disease forecasting, and several important questions remain unresolved. First, the optimal neural architecture for this task is unclear. While our hybrid CNN-transformer design performs well, alternative architectures may offer better performance or efficiency trade-offs. Second, the ideal balance between simulation diversity and computational cost requires further investigation. Our training set spans 400 million simulated days, generated continuously until model convergence without repeating data. Using synthetic data like this allows us to avoid the overfitting risks inherent in epoch-based training on fixed datasets. However, whether this scale is necessary or whether comparable performance could be achieved with fewer but more strategically designed simulations remains to be seen.
@@
Mantis illustrates a path forward for general-purpose disease forecasting systems that can be quickly deployed without disease-specific tuning. Its out-of-the-box generalization makes it especially well-suited for early outbreak response, rural regions, and low- and middle-income countries—settings where traditional models often fail due to sparse or delayed data. By pairing simulation-based priors with flexible neural architectures, this approach offers a scalable alternative to expert-tuned, disease-specific models. As a first demonstration of simulation-grounded foundation models in infectious disease forecasting, Mantis establishes both the promise and the open challenges of this paradigm for future decision support systems in global public health.
@@ Methods
CD led the study and developed the code and initial draft of the manuscript. CH and AS did the ablation studies and RM developed figure visualizations. CD, MCE, RM, and EM all contributed substantial edits and revisions for the final manuscript.
@@ Appendix A Training Data
This comprehensive parameter space enables the generation of diverse outbreak scenarios spanning different pathogen characteristics, population contexts, surveillance qualities, and intervention regimes. Each simulation run samples independently from these distributions, creating a rich training corpus that captures the full spectrum of human-to-human transmission dynamics.
@@ Appendix D Additional Results
D.1 Ablation Studies Confirm the Importance of Scale and Mechanistic Pretraining
To understand the key factors driving Mantis’s performance, we conducted two ablation studies examining the roles of architectural capacity and mechanistic pretraining in enabling accurate plug-and-play forecasting.
D.1.1 Architectural Capacity Requirements
Despite identical training data and procedures, the reduced architectural capacity led to substantially degraded performance. When evaluated on COVID-19 mortality forecasting across U.S. states, the LSTM architecture achieved mean absolute error (MAE) that was 89% higher than the full Mantis model. This demonstrates that while the synthetic training paradigm is powerful, it requires sufficient architectural capacity to effectively learn and generalize from the diverse mechanistic patterns encoded in the simulation data.
D.1.2 Mechanistic Pretraining is Essential for Novel Outbreak Performance
To isolate the contribution of mechanistic pretraining, we trained the full Mantis architecture exclusively on real-world data, eliminating simulations from the training process. We evaluated on COVID-19 hospitalization forecasting. We implemented a rolling training strategy that provided the strongest possible baseline: the model was initially trained on the first few months of pandemic data, then progressively retrained with expanding datasets as more real-world observations became available. This approach mimics operational practice and ensures that the real-world-only model has access to all available historical information at each forecast point.
The results reveal the critical importance of mechanistic pretraining. When trained only on real-world COVID-19 data, forecast error increased by 262% compared to the simulation-pretrained Mantis model. This dramatic performance degradation reflects the fundamental challenge of learning outbreak dynamics from limited historical data during a novel pandemic. Real-world training data, even when accumulated over months, cannot capture the full spectrum of epidemic behaviors that might emerge. In contrast, the simulation-pretrained model draws from hundreds of millions of days of diverse outbreak scenarios, enabling it to recognize and extrapolate patterns even in unprecedented contexts.
These ablation results underscore the core insight behind Mantis: mechanistic pretraining enables models to internalize generalizable principles of epidemic dynamics that extend far beyond what can be learned from any single real-world outbreak. While architectural sophistication contributes to performance, the simulation-grounded training paradigm is the essential component that enables out-of-the-box generalization to novel diseases and outbreak contexts.
D.2 Performance Consistency Across Population Scales

### v4

@@ Discussion
Our results using Mantis demonstrate that accurate forecasting across diverse diseases can be achieved through training on simulated outbreaks, without access to real-world data. By learning from mechanistic simulations, Mantis delivers competitive performance across real-world forecasting tasks spanning multiple transmission modes, outbreak phases, and surveillance systems. These capabilities position Mantis and other simulation-based foundation models as a practical tool for public health use, especially in low-resource or emerging outbreak settings where both historical data and specialized modeling expertise are limited. Because Mantis can generate forecasts with only a short initial observation window, it enables early situational awareness when conventional models require longer histories to achieve comparable accuracy. Mantis’s advantage in data-scarce settings stems from its ability to avoid overfitting to sparse initial data. Rather than extrapolating from a handful of observations, Mantis produces appropriately hedged forecasts that reflect the true uncertainty across potential outbreak trajectories in our evaluations, enabling more reliable probabilistic planning even when historical context is minimal. Mantis avoids overfitting because using synthetic data allows us to generate as many simulations as necessary, avoiding the overfitting or memorization risks inherent in epoch-based training on fixed datasets.
@@
This approach offers a potential solution to a fundamental challenge in epidemiological modeling: real-world disease data are limited by data availability, reporting inconsistencies, and the infrequency of major outbreaks. Simulations, by contrast, can generate unlimited training data spanning diverse scenarios. Whether simulation-based training ultimately proves superior to training on massive real-world datasets—if such data could be assembled—remains an open question. However, given current data availability constraints, the simulation-based approach offers a practical path for building general-purpose forecasting systems.
It is important to note, though, that this approach depends critically on developing simulations that are sufficiently realistic and diverse to capture the dynamics a model will encounter in practice. Our results suggest that a broad distribution of acute outbreak simulations can generalize even to chronic disease dynamics not explicitly modeled, but the boundaries of this generalization are unclear. Future work should investigate when simulation-trained models fail and whether certain epidemic regimes require explicit representation in training, or whether sufficiently diverse simulations can approximate arbitrary contagion dynamics.
@@
Mantis represents an initial attempt at a foundation model for disease forecasting, and several important questions remain unresolved. First, the optimal neural architecture for this task is unclear. While our hybrid CNN-transformer design performs well, alternative architectures may offer better performance or efficiency trade-offs. Second, the ideal balance between simulation diversity and computational cost requires further investigation. Our training set spans 400 million simulated days, generated continuously until model convergence without repeating data. Whether this scale is necessary or whether comparable performance could be achieved with fewer but more strategically designed simulations remains to be seen. More real-time benchmarking of Mantis will also more clearly demonstrate the settings where it succeeds and where it struggles, which will become more clear in coming seasons.
@@
Mantis illustrates a potential path forward for general-purpose disease forecasting systems that can be quickly deployed without disease-specific tuning. Its out-of-the-box generalization makes it especially well-suited for early outbreak response, rural regions, and low- and middle-income countries—settings where traditional models often fail due to sparse or delayed data. By pairing simulation-based priors with flexible neural architectures, this approach offers a scalable alternative to expert-tuned, disease-specific models. As a first demonstration of simulation-grounded foundation models in infectious disease forecasting, Mantis establishes both the promise and the open challenges of this paradigm for future decision support systems in global public health.
@@ Methods
CD led the study and developed the code and initial draft of the manuscript. CH and AS did the ablation studies and RM developed figure visualizations. CD, MCE, and RM all contributed substantial edits and revisions for the final manuscript.
Acknowledgments
This project was made possible by the Insight Net cooperative agreement with University of Michigan (5 NU38FT000002-02-00) from the CDC’s Center for Forecasting and Outbreak Analytics (CDC-RFA-FT-23-0069). Its contents are solely the responsibility of the authors and do not necessarily represent the official views of the Centers for Disease Control and Prevention. The authors thank Julie Gilbert for her assistance with gathering data for an earlier version of the manuscript.
@@ Appendix A Training Data
This large, high dimensional parameter space enables the generation of diverse outbreak scenarios spanning different pathogen characteristics, population contexts, surveillance qualities, and intervention regimes. Each simulation run samples independently from these distributions, creating a rich training corpus that captures the full spectrum of human-to-human transmission dynamics.
@@ Appendix D Additional Results
D.1 Performance Consistency Across Population Scales